# Enhancing Photon Identification with Neural Network Methods

### v1

@@ Abstract
We investigate photon–pion discrimination in regimes where electromagnetic showers overlap at the scale of calorimeter granularity. Using full detector simulations with fine-grained calorimeter segmentation of approximately $0.025\times 0.025$ in $(\eta,\phi)$, we benchmark three approaches: boosted decision trees (BDTs) on shower-shape variables, dense neural networks (DNNs) on the same features, and a ResNet-based convolutional neural network operating directly on calorimeter cell energies. The ResNet significantly outperformed both baseline methods, achieving further gains when augmented with soft scoring and an auxiliary $\Delta R$ regression head. Our results demonstrate that residual convolutional architectures, combined with physics-informed loss functions, can substantially improve photon identification in high-luminosity collider environments in which overlapping electromagnetic showers challenge traditional methods.
@@ 1 Introduction
The Large Hadron Collider (LHC) has enabled precision measurements of Standard Model processes and opened new frontiers in the search for physics beyond the Standard Model. Central to many of these studies is the accurate identification of photons, which provide relatively clean experimental signatures in the high-multiplicity environment of hadron collisions. The clean signature of photons in hadron colliders has made them a powerful probe for precision measurements and searches for new physics [ATLAS:2012gk, CMS:2012qbp]. Therefore, accurate reconstruction and identification of photons is a crucial task in high-energy physics experiments [ATLAS:2018egammaPerf]. Photon candidates are identified and reconstructed using calorimeter and tracker information. Due to the nature of photon interactions in the calorimeter – namely, pair production and bremsstrahlung – a photon entering the detector produces a cascade of secondary particles, forming an electromagnetic (EM) shower [Leo:1994Introduction].
@@
Traditionally, EM showers are characterized using a set of carefully engineered observables known as shower-shape variables [ATLAS:2019PhotonID]. Selection cuts or multivariate classifiers trained on these variables are employed to distinguish prompt photons, produced directly in the hard scattering process, from non-prompt photons originating from hadronic decays or fragmentation. In both the ATLAS and CMS experiments, boosted decision trees (BDTs) based on these shower-shape variables constitute the standard photon-identification algorithm [ATLAS:2018egammaPerf, ATLAS:2019PhotonID, CMS]. These approaches have demonstrated excellent performance across most kinematic regimes, but their discriminating power degrades in challenging conditions such as high pile-up (multiple overlapping proton–proton collisions in the same bunch crossing) or when EM showers become spatially overlapping at fine calorimeter granularity. In recent years, machine learning techniques have increasingly been explored for photon and electron identification, aiming to surpass the limitations of handcrafted shower-shape variables. Early studies by the ATLAS and CMS collaborations demonstrated that deep neural networks trained on calorimeter observables can improve discrimination power compared to traditional BDT-based approaches [ATLAS_CNNPhoton, CMS_DNNPhoton]. Further work has extended these ideas to more expressive architectures such as convolutional neural networks (CNNs) operating directly on calorimeter cell images [deOliveira2016, Accettura2023], or hybrid models combining high-level variables with low-level features [Duarte2018, Komiske2019]. Collectively, these studies highlight the growing potential of deep learning in EM shower classification tasks. However, despite these advances, a persistent challenge in photon identification remains: the discrimination of neutral pion ($\pi^{0}$) decays that mimic single-photon signatures.
@@
A notable difficulty arises from the presence of $\pi^{0}$ mesons, which decay to two photons in over $98\%$ of cases [PDG2024]. For a $\pi^{0}$ mass of ${\sim}135\,\text{MeV}$ and photon energies in the GeV range, the opening angle $\theta$ between the two decay photons is typically very small [ATLAS:2018egammaPerf, PDG2024]. This understanding follows from the invariant mass relation for massless photons:
@@
which shows that higher-energy photons correspond to smaller opening angles. As a result, the two decay photons can be highly collimated, sometimes depositing their energy in overlapping regions of the calorimeter, making them difficult to distinguish from single photons (see Fig. 1) [ATLAS:2018egammaPerf].
@@
In this work, we investigate the use of deep neural networks for photon identification in a challenging kinematic regime where standard techniques exhibit reduced performance: photon energies between 10 and 100 GeV, where collimated $\pi^{0}$ decays are most problematic. Simulated samples are generated with the COCOA-HEP library [DiBello:2023cocoa], which couples PYTHIA8 [Sjostrand:2014zea] event generation with full detector simulation in Geant4 [Agostinelli:2002hh]. The calorimeter was configured with three layers in both the electromagnetic and hadronic sections. The EM1 and EM2 layers have nearly symmetric granularity of approximately $0.025\times 0.025$ in $(\eta,\phi)$, with exact segmentation shown in Table 1. This near-uniform segmentation provides a suitable two-dimensional shower representation for convolutional neural network studies.
We benchmark three complementary approaches: (i) the standard shower-shape variables with a BDT baseline [ATLAS:2018egammaPerf], (ii) a dense neural network trained on the same high-level inputs, and (iii) a ResNet-based convolutional network trained directly on calorimeter images. Additionally, we explore the impact of separating background events according to their diphoton separation $\Delta R$, investigating whether dedicated training for events with $\Delta R<0.025$ yields further performance gains. While earlier studies have primarily addressed photon–electron or photon–jet separation using deep learning, the present work focuses specifically on the challenging case of $\pi^{0}\to\gamma\gamma$ overlaps at high calorimeter granularity. In addition, we introduce two physics-informed strategies – soft scoring and an auxiliary $\Delta R$ regression head—that guide the network toward spatially discriminative features beyond those employed in previous approaches.
@@ 2 Monte Carlo Sample Generation
For this study, a calorimeter simulation with high transverse and longitudinal granularity was utilized to perform photon identification (PID) using a CNN on detailed shower-shape information. While fast-simulation frameworks such as DELPHES [DELPHES] can be configured to achieve a transverse segmentation comparable to the one used here (down to $\eta\times\phi$ cells of $0.025\times 0.025$ or finer), they are inherently limited to a single EM and a single hadronic calorimeter layer [ATL-PHYS-PUB-2017-015, CMS:2017dal]. These layers are modeled as 2D energy maps derived from parameterized detector resolutions, lacking a realistic longitudinal shower development. As a result, multi-layer correlations, depth-dependent energy patterns, and realistic shower-shapes cannot be reproduced. Such information is essential for CNN-based PID, making DELPHES inadequate for the purposes of this study.
Instead, we utilized the COCOA-HEP detector simulation [DiBello:2023cocoa], which provides full 3D calorimeter modeling with multiple configurable layers, per-cell energy deposits, and realistic shower evolution. The collision and detector parameters were chosen to be broadly consistent with ATLAS in terms of magnetic field, radiation length, and $\sqrt{s}=14$ TeV, corresponding to the design energy of the High-Luminosity LHC phase in which future photon identification studies will operate. The calorimeter was configured with three layers in both the EM and hadronic sections, each with uniform $\eta\times\phi$ segmentation as shown in Table 1.
@@ 2.1 Signal Samples
Signal events were generated from five prompt-photon production processes: $g\rightarrow q\gamma, qq$$\rightarrow g\gamma, qq$$q\rightarrow\gamma\gamma$, $gg\rightarrow g\gamma$, and $gg\rightarrow\gamma\gamma$ [Sjostrand:2014zea]. A minimum hard-scatter transverse momentum (pTHatMin) of 10 GeV was applied, and only photons with $10<p_{T}<100$ GeV were retained. The energy range between 10 and 100 GeV was selected as this is the most challenging regime for photon identification. Above 100 GeV, identification efficiency saturates at a high plateau[ATLAS:2018egammaPerf]; while below 10 GeV, EM showers are often too small and poorly contained to be reliably separated from hadronic backgrounds, making photon identification practically impossible under realistic detector conditions [ATLAS:2018egammaPerf, ATLAS:2019PhotonID]. From the full generated dataset, 1.1 million events were selected to produce a flat $p_{T}$ spectrum, avoiding the over-representation of low-$p_{T}$ photons inherent in the original steeply falling distribution. The dataset was then split into 70% for training and validation, and 30% for testing.
@@ 2.2 Background Samples
Background events were generated from all QCD processes, requiring at least one $\pi^{0}$ in the final state with $10<p_{T}<100$ GeV, decaying into two photons [Sjostrand:2014zea, PDG2024]. distribution of $\Delta R$ between the two photons is shown in figure 4, with two thirds of the diphoton pairs having $\Delta R<0.025$, lower than the detector’s resolution. Unlike the signal, no $p_{T}$ reweighting was performed on the background, preserving the true steeply falling QCD spectrum in which low-$p_{T}$ events dominate.
A minimum hard-scatter transverse momentum (pTHatMin) of 10 GeV was applied [Sjostrand:2014zea]. Unlike the signal, no $p_{T}$ reweighting was performed on the background. This decision preserves the true steeply falling QCD $p_{T}$ spectrum, which is essential because the background classifier must be trained to reject events across the realistic, dominant low-$p_{T}$ phase space observed in experimental data.
@@ 3.1 Shower-Shape and BDT Baseline
Algorithms currently used for object classification in collider experiments are often referred to as Cut-Based algorithms. For each object, a set of shower-shape variables is calculated, and a multi-dimensional cut is optimized to select signal objects at a given fake-positive rate. In practice, modern implementations frequently use a Boosted Decision Tree (BDT) trained on the same variables instead of a literal cut-flow. In this work, we implement the BDT as baseline, using XGBoost [Chen:2016XGBoost]. The 20 shower-shape variables used in this work are based on those employed in ATLAS photon identification analyses [ATLAS:2019PhotonID] and are listed in Table 3.
@@ 3.2 Dense Neural Network Baseline
To provide a direct comparison with the BDT, we also implement a Dense Neural Network (DNN) trained on the same twenty shower-shape variables. This baseline, implemented in PyTorch [Paszke:2019PyTorch], is designed to have a comparable number of trainable parameters to the BDT, ensuring that differences in performance are due to model architecture rather than capacity. The DNN consists of several fully connected layers with ReLU activations and dropout for regularization, as shown in 9.
@@ 3.3 CNN/ResNet Calorimeter-Based Model
Each branch is processed through a sequence of residual convolutional blocks (ResNet blocks), allowing the network to learn deep spatial features without vanishing gradients. The outputs of the EM and Had branches are flattened, concatenated, and passed through fully connected layers for classification, as shown in 10. This two-prong ResNet-based architecture, implemented in PyTorch [Paszke:2019PyTorch] and using residual blocks [He:2015ResNet], is designed to exploit spatial correlations in the calorimeter data, which are inaccessible to cut-based or shower-shape approaches.
@@ 3.4 Additional Training Strategies
Beyond the baseline ResNet architecture, we explored two complementary strategies aimed at improving performance in challenging background regimes.
@@
Background events with photon pairs at very small angular separation ($\Delta R<0.025$) are particularly difficult to reject due to their high similarity to true photon showers. To encourage the network to better separate these cases, we implemented a soft scoring scheme (also known as label smoothing) [Szegedy:2016SoftLabels, müller:smoothing] in which such background events are assigned continuous target labels between 0 and 1 rather than a hard zero, as shown in 4. This smooth target transition reduces the loss penalty for near-misses and allows the model to focus its capacity on improving discrimination where it is most needed. Following multiple tests, the Fermi-Dirac (FD) function in the range [0,0.4] achieved the best results.
Figure 4: Distribution of $\Delta R$ between photon pairs from $\pi^{0}$ decay (blue histogram, left axis) and the corresponding soft-scoring target function (red curve, right axis). Background events ware assigned continuous labels between 0 and 0.4 using a Fermi-Dirac-inspired function, reducing the penalty for near-threshold cases while maintaining hard labels for well-separated pairs.
@@
We further added an auxiliary output branch that predicts the $\Delta R$ of background photon pairs [Caruana:1997Multitask], as shown in 11. This branch shares the same convolutional backbone as the main classifier but has its own regression head, contributing an auxiliary loss term during training. The predicted $\Delta R$ does not contribute to the final classification score, as it is highly correlated with the output score. Rather, it serves as an additional supervision signal to guide the feature extractor towards representations sensitive to small-scale spatial differences. This multi-task setup proved effective in regularizing the network and improving robustness, especially for closely spaced photon pairs.
@@ 4 Results
This section presents the performance of the algorithms introduced in Section 3. We first compare the baseline approaches — shower-shape-based BDT, a DNN with the same inputs, and a calorimeter-based ResNet — to quantify the gain from using full calorimeter information. We then evaluate refinements to the ResNet, including the soft-scoring approach and the auxiliary $\Delta R$ head. Finally, performance is shown separately for two background topologies, defined by $\Delta R<0.025$ and $\Delta R\geq 0.025$ between the two photons from $\pi^{0}$ decay.
4.1 Baseline performance: BDT, DNN, and ResNet

### v2

@@ Abstract
We investigate photon–pion discrimination in regimes where electromagnetic showers overlap at the scale of calorimeter granularity. Using full detector simulations with fine-grained calorimeter segmentation of approximately $0.025\times 0.025$ in $(\eta,\phi)$, we benchmark three approaches: boosted decision trees (BDTs) on shower-shape variables, dense neural networks (DNNs) on the same features, and a ResNet-based convolutional neural network operating directly on calorimeter cell energies. The ResNet significantly outperformed both baseline methods, achieving further gains when augmented with soft scoring and an auxiliary $\Delta R$ regression head. Our results demonstrate that residual convolutional architectures, combined with physics-motivated training strategies, can substantially improve photon identification in high-luminosity collider environments in which overlapping electromagnetic showers challenge traditional methods.
@@ 1 Introduction
The Large Hadron Collider (LHC) has enabled precision measurements of Standard Model processes and opened new frontiers in the search for physics beyond the Standard Model. Central to many of these studies is the accurate identification of photons, which provide relatively clean experimental signatures in the high-multiplicity environment of hadron collisions. The clean signature of photons in hadron colliders has made them a powerful probe for precision measurements and searches for new physics [4, 10]. Therefore, accurate reconstruction and identification of photons is a crucial task in high-energy physics experiments [5]. Photon candidates are identified and reconstructed using calorimeter and tracker information. Due to the nature of photon interactions in the calorimeter – namely, pair production and bremsstrahlung – a photon entering the detector produces a cascade of secondary particles, forming an electromagnetic (EM) shower [18].
@@
Traditionally, EM showers are characterized using a set of carefully engineered observables known as shower-shape variables [6]. Selection cuts or multivariate classifiers trained on these variables are employed to distinguish prompt photons, produced directly in the hard scattering process, from non-prompt photons originating from hadronic decays or fragmentation. In both the ATLAS and CMS experiments, boosted decision trees (BDTs) based on these shower-shape variables constitute the standard photon-identification algorithm [5, 6, 12]. These approaches have demonstrated excellent performance across most kinematic regimes, but their discriminating power degrades in challenging conditions such as high pile-up (multiple overlapping proton–proton collisions in the same bunch crossing) or when EM showers become spatially overlapping at fine calorimeter granularity. In recent years, machine learning techniques have increasingly been explored for photon and electron identification, aiming to surpass the limitations of handcrafted shower-shape variables. Early studies by the ATLAS and CMS collaborations demonstrated that deep neural networks trained on calorimeter observables can improve discrimination power compared to traditional BDT-based approaches [22, 21]. Further work has extended these ideas to more expressive architectures such as convolutional neural networks (CNNs) operating directly on calorimeter cell images [14, 1], or hybrid models combining high-level variables with low-level features [15, 17]. Collectively, these studies highlight the growing potential of deep learning in EM shower classification tasks. However, despite these advances, a persistent challenge in photon identification remains: the discrimination of neutral pion ($\pi^{0}$) decays that mimic single-photon signatures.
@@
A notable difficulty arises from the presence of $\pi^{0}$ mesons, which decay to two photons in over $98\%$ of cases [19]. For a $\pi^{0}$ mass of ${\sim}135\,\text{MeV}$ and photon energies in the GeV range, the opening angle $\theta$ between the two decay photons is typically very small [5, 19]. This understanding follows from the invariant mass relation for massless photons:
@@
which shows that higher-energy photons correspond to smaller opening angles. As a result, the two decay photons can be highly collimated, sometimes depositing their energy in overlapping regions of the calorimeter, making them difficult to distinguish from single photons (see Fig. 1) [5].
@@
In this work, we investigate the use of deep neural networks for photon identification in a challenging kinematic regime where standard techniques exhibit reduced performance: photon energies between 10 and 100 GeV, where collimated $\pi^{0}$ decays are most problematic. Simulated samples are generated with the COCOA-HEP library [7], which couples PYTHIA8 [23] event generation with full detector simulation in Geant4 [2]. The calorimeter was configured with three layers in both the electromagnetic and hadronic sections. The EM1 and EM2 layers have nearly symmetric granularity of approximately $0.025\times 0.025$ in $(\eta,\phi)$, with exact segmentation shown in Table 1. This near-uniform segmentation provides a suitable two-dimensional shower representation for convolutional neural network studies.
We benchmark three complementary approaches: (i) the standard shower-shape variables with a BDT baseline [5], (ii) a dense neural network trained on the same high-level inputs, and (iii) a ResNet-based convolutional network trained directly on calorimeter images. Additionally, we explore the impact of separating background events according to their diphoton separation $\Delta R$, investigating whether dedicated training for events with $\Delta R<0.025$ yields further performance gains. While earlier studies have primarily addressed photon–electron or photon–jet separation using deep learning, the present work focuses specifically on the challenging case of $\pi^{0}\to\gamma\gamma$ overlaps at high calorimeter granularity. In addition, we introduce two physics-motivated strategies – soft scoring and an auxiliary $\Delta R$ regression head—that guide the network toward spatially discriminative features beyond those employed in previous approaches.
@@ 2 Monte Carlo Sample Generation
For this study, a calorimeter simulation with high transverse and longitudinal granularity was utilized to perform photon identification (PID) on detailed shower-shape information. While fast-simulation frameworks such as DELPHES [13] can be configured to achieve a transverse segmentation comparable to the one used here (down to $\eta\times\phi$ cells of $0.025\times 0.025$ or finer), they are inherently limited to a single EM and a single hadronic calorimeter layer [3, 11]. These layers are modeled as 2D energy maps derived from parameterized detector resolutions, lacking a realistic longitudinal shower development. As a result, multi-layer correlations, depth-dependent energy patterns, and realistic shower-shapes cannot be reproduced. Such information is essential for accurate PID, making DELPHES inadequate for the purposes of this study.
Instead, we utilized the COCOA-HEP detector simulation [7], which provides full 3D calorimeter modeling with multiple configurable layers, per-cell energy deposits, and realistic shower evolution. The collision and detector parameters were chosen to be broadly consistent with ATLAS in terms of magnetic field, radiation length, and $\sqrt{s}=14$ TeV, corresponding to the design energy of the High-Luminosity LHC phase. The calorimeter was configured with three layers in both the EM and hadronic sections, each with uniform $\eta\times\phi$ segmentation as shown in Table 1.
To maximize the realism of the simulation, electronic noise was implemented via the COCOA framework using layer-dependent Gaussian noise: [13, 34, 41] MeV for the EM layers and [75, 50, 25] MeV for the hadronic layers. These values are based on the best imitation of the ATLAS detector environment, ensuring that the simulated energy deposits include the stochastic fluctuations present in experimental data.
@@ 2.1 Signal Samples
Signal events were generated from five prompt-photon production processes: $g\rightarrow q\gamma, qq$$\rightarrow g\gamma, qq$$q\rightarrow\gamma\gamma$, $gg\rightarrow g\gamma$, and $gg\rightarrow\gamma\gamma$ [23]. A minimum hard-scatter transverse momentum (pTHatMin) of 10 GeV was applied, and only photons with $10<p_{T}<100$ GeV were retained. The energy range between 10 and 100 GeV was selected as this is the most challenging regime for photon identification. Above 100 GeV, identification efficiency saturates at a high plateau [5]; while below 10 GeV, EM showers are often too small and poorly contained to be reliably separated from hadronic backgrounds, making photon identification practically impossible under realistic detector conditions [5, 6]. From the full generated dataset, 1.1 million events were selected to produce a flat $p_{T}$ spectrum, avoiding the over-representation of low-$p_{T}$ photons inherent in the original steeply falling distribution. The dataset was then split into 70% for training and validation, and 30% for testing.
@@ 2.2 Background Samples
Background events were generated from all QCD processes, requiring at least one $\pi^{0}$ in the final state with $10<p_{T}<100$ GeV, decaying into two photons [23, 19]. The distribution of $\Delta R$ between the two photons is shown in Figure 4, with two-thirds of the diphoton pairs having $\Delta R<0.025$, lower than the detector’s resolution.
A minimum hard-scatter transverse momentum (pTHatMin) of 10 GeV was applied [23]. Unlike the signal, no $p_{T}$ reweighting was performed on the background. This decision preserves the true steeply falling QCD $p_{T}$ spectrum, which is essential to reflect the dominant low-$p_{T}$ phase space observed in experimental data.
@@ 3.1 Shower-Shape and BDT Baseline
Algorithms currently used for object classification in collider experiments are often referred to as Cut-Based algorithms. For each object, a set of shower-shape variables is calculated, and a multi-dimensional cut is optimized to select signal objects at a given fake-positive rate. In practice, modern implementations frequently use a Boosted Decision Tree (BDT) trained on the same variables instead of a literal cut-flow. In this work, we implement the BDT as baseline, using XGBoost [9]. The 20 shower-shape variables used in this work are based on those employed in ATLAS photon identification analyses [6] and are listed in Table 3.
@@ 3.2 Dense Neural Network Baseline
To provide a direct comparison with the BDT, we also implement a Dense Neural Network (DNN) trained on the same twenty shower-shape variables. This baseline, implemented in PyTorch [20], is designed to have a comparable number of trainable parameters to the BDT, ensuring that differences in performance are due to model architecture rather than capacity. The DNN consists of several fully connected layers with ReLU activations and dropout for regularization, as shown in Figure 11.
@@ 3.3 CNN/ResNet Calorimeter-Based Model
Each branch is processed through a sequence of residual convolutional blocks (ResNet blocks), allowing the network to learn deep spatial features without vanishing gradients. The outputs of the EM and Had branches are flattened, concatenated, and passed through fully connected layers for classification, as shown in Figure 12. This two-prong ResNet-based architecture, implemented in PyTorch [20] and using residual blocks [16], is designed to exploit spatial correlations in the calorimeter data, which are inaccessible to cut-based or shower-shape approaches.
@@ 3.4 Additional Training Strategies
We explore two physics-motivated training strategies designed to address specific challenges in the overlapping shower regime.
@@
Background events with photon pairs at very small angular separation ($\Delta R<0.025$) are particularly difficult to reject due to their high similarity to true photon showers. To address label ambiguity in this regime, we implemented a soft scoring scheme (also known as label smoothing) [24, müller:smoothing] in which such background events are assigned continuous target labels between 0 and 1 rather than a hard zero, as shown in Figure 4. This smooth target transition reduces the loss penalty for near-misses and allows the model to focus its capacity on improving discrimination where it is most needed. We systematically evaluate three functional forms—Fermi-Dirac (FD), linear, and exponential—across a range of maximum score amplitudes (0.1–0.6), with detailed results presented in Section 4.2.
Figure 4: Distribution of $\Delta R$ between photon pairs from $\pi^{0}$ decay (histogram, left axis) and three soft-scoring target functions (right axis). We evaluate Fermi-Dirac (FD), linear, and exponential functional forms, each shown for a representative maximum score amplitude. Background events are assigned continuous labels between 0 and the maximum amplitude using these functions, reducing the penalty for near-threshold cases while maintaining hard labels for well-separated pairs.
@@
We further added an auxiliary output branch that predicts the $\Delta R$ of background photon pairs [8], as shown in Figure 13. This branch shares the same convolutional backbone as the main classifier but has its own regression head, contributing an auxiliary loss term during training. The predicted $\Delta R$ does not contribute to the final classification score. Rather, it serves as an regularization mechanism, encouraging the shared feature extractor to learn representations sensitive to spatial shower structure. The auxiliary loss weight $\alpha$ controls the relative contribution of classification (BCE) versus regression (MSE) objectives, with systematic evaluation presented in Section 4.3. This multi-task setup proved effective in regularizing the network and improving robustness, especially for closely spaced photon pairs.
3.5 Experimental Protocol
To ensure fair comparison across all experiments, we employ controlled initialization and data ordering. All models in the ablation studies (Sections 4.2, 4.3, and 4.4) are initialized from identical random weights (seed=42) and trained on data presented in the same order across epochs. This eliminates initialization variance and ensures that performance differences reflect architectural choices or training strategies rather than stochastic effects. Each configuration is trained for 100 epochs using the Adam optimizer with learning rate $10^{-3}$ and batch size 1000.
@@ 4 Results
This section presents comprehensive performance evaluation of the proposed methods, including systematic hyperparameter optimization, ablation studies, and robustness tests under detector systematics.
We first compare baseline approaches to quantify the gain from using full calorimeter information over engineered shower-shape variables (Section 4.1). We then systematically optimize two physics-motivated training strategies: soft scoring to address label ambiguity (Section 4.2) and auxiliary $\Delta R$ supervision for feature regularization (Section 4.3). Section 4.4 evaluates the combination of these techniques, revealing interference effects when using independently-optimized hyperparameters. Finally, Section 4.5 assesses robustness under realistic detector systematics including energy scale variations and increased noise conditions.
4.1 Baseline Comparison: BDT, DNN, and ResNet
This subsection establishes baseline performance for three approaches: shower-shape variables with a BDT, a dense neural network (DNN) using the same inputs, and a ResNet operating directly on calorimeter images. All performance curves include binomial uncertainty bands.